Show that
[ x,e iap ] = -ae iap
h h
1
Expert's answer
2020-04-13T10:17:07-0400

In order to evaluate commutator "[x, e^{i a p}]", we need to expand the exponent into Taylor series, i.e. "e^{i a p} = 1 + i a p + \frac{(i a p)^2}{2!} + \frac{(i a p)^3}{3!} + ...".


Since the commutator is linear in both arguments, we need to evaluate commutators like "[x, p^n]". Using Leibniz rule for commutators and canonical commutation relation "[x, p] = i \hbar":


"[x, p^n] = (i \hbar) p^{n-1} + p (i \hbar) p^{n-2} + ... + p^{n-1} (i \hbar) = i \hbar n p^{n-1}".


Hence,

"[x, e^{i a p}] = [x, 1 + i a p + \frac{(i a p)^2}{2!} + \frac{(i a p)^3}{3!} + ...] ="

"= (i \hbar)[(i a)^1 + \frac{(i a)^2}{2!} \cdot 2 \cdot p^1 + \frac{(i a)^3}{3!} \cdot 3 \cdot p^2 +... = (i \hbar) (i a) \sum_{n=0}^{\infty} \frac{(i a p)^n}{n!} = - a \hbar e^{i a p}".